A wooden incline plane is being built, along which wooden crates of merchandise are to be slid down into the basement of a store.

a.what angle with the horizontal should the incline plane be made if the crates are to slide down at constant speed?

b.With what force must a 100 kg crate be pushed in order to start it sliding down the incline plane if the angle of the incline plane is that the one found in (a)?
1
Expert's answer
2020-12-06T17:25:14-0500


a. For crates sliding down at constant speed along the x axis (see figure), the friction force F should be balanced by the projection of the gravity force on the x axis, that equal "mgsin\alpha".

Friction force is "F=\mu N",

where "N=mgcos\alpha" (reaction of the incline plane),

"\mu" - coefficient of sliding friction, wich for sliding of wood on wood is 0.48.

"mgsin\alpha=\mu mgcos\alpha,"

"tan\alpha=\mu",

"\alpha=arctan\mu = arctan0.48=26 \space degrees."


b. To start a 100 kg crate sliding down the incline plane the pushing force should reach the value of static friction force, that is "\mu_sN,"

where "\mu_s" - coefficient of static friction, wich for wood on wood is 0.62.

So,

"F_p=\mu_s mgcos\alpha=0.62\cdot100\cdot9.81\cdot \cos26\degree=546.7 N"

Answer: a: 26 deg, b: 546.7 N.